OpenStudy (anonymous):

convert each rectangular equation into polar form x^2+y^2=3

OpenStudy (anonymous):

that sounds really complicated

OpenStudy (anonymous):

srry i cant help

myininaya (myininaya):

x=rcos(theta) y=rsin(theta)

myininaya (myininaya):

that should help

OpenStudy (anonymous):

Or, x^2+y^2=r^2

OpenStudy (anonymous):

r^2 = x^2 + y^2 theta = arctan(y/x)